The number of carbohydrates found in a random sample of fast-food entrees is listed below. Is there sufficient evidence to conclude that the variance differs from 100? use the 0.05 level of significance.


53 46 39 39 30

47 38 73 43 41


1
Expert's answer
2021-07-12T18:39:44-0400

x= number of carbohydrates

"n=10 \\

\bar{x}= \frac{53+46+...+43+41}{10}=44.9 \\

S^2 = \frac{1}{n-1} \sum (x_i - \bar{x})^2 \\

S^2 = \frac{1}{10-1} \times ((53-44.9)^2 + (46-44.9)^2+...+(43-44.9)^2 + (41-44.9)^2) = 135.53"

We have to test:

"H_0: \sigma^2=100 \\

H_1: \sigma^2 ≠ 100 \\

α=0.05"

Test-statistic:

"χ^2 = \frac{(n-1)S^2}{\sigma^2} \\

= \frac{9 \times 135.43}{100} = 12.188"

Critical value "\u03c7^2_{n-1, \u03b1\/2}=19.023"

"\u03c7^2" < Critical value

Accept H0.

There is not sufficient evidence to conclude that the variance differs from 100.